2/3g+1/2g=14
I don't have a way to write fractions, but the above / is a fraction. I do not know how to solve it. Please help :)

Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

2/3g+1/2g=14
I don't have a way to write fractions, but the above / is a fraction. I do not know how to solve it.
-------------------
That's how fractions are written.
Be careful to show it's (2/3)g and not 2/(3g), tho.
----------
(2/3)g + (1/2)g=14
Add the 2 g terms.
Find a common denominator to do that.
The product always works.
2*3 = 6
-----------
(2/3)g + (1/2)g=14
(4/6)g + (3/6)g=14
(7/6)g = 14
Multiply both sides by 6
7g = 84
Divide by 7
g = 12

Answer by stanbon(75887)

2/3g+1/2g=14
-----
Mltiply thru by 6 to get:
2*2g + 3g = 84
7g = 84
g = 12
==============
Cheers,
Stan H.